OpenStudy (anonymous):

add and simplify 9y/y^2 + y/y-8

jhonyy9 (jhonyy9):

\[9y/y ^{2}=9/y\]

jhonyy9 (jhonyy9):

9/y +y/(y-8)=((9(y-8)+y(y))/y(y-8)=(9y-72+y(y))/y(y-8)

jhonyy9 (jhonyy9):

\[(y ^{2}+9y-72)/y(y-8)\]